### v2

# Highly-Efficient Differentiable Simulation for Robotics

Abstract
Over the past few years, robotics simulators have significantly improved in computational speed and scalability, enabling them to generate years of simulated data for complex robotic systems in a matter of minutes or hours. However, despite these advancements, the efficient and accurate computation of simulation derivatives remains an open challenge. Addressing this challenge could substantially accelerate the convergence of reinforcement learning and trajectory optimization algorithms, particularly for contact-rich problems. This paper addresses this challenge by introducing a unifying framework for robotic simulation that comprehensively considers all the factors in simulation, including dynamics, collisions, and friction. It results in an efficient algorithm that leverages implicit differentiation to compute the analytical derivatives of the simulation. It explicitly accounts for the intrinsic non-smoothness of collision and frictional simulation stages while exploiting the sparsity in dynamics induced by the multi-body system structure. These derivatives have been implemented in C++, and the code will be open-sourced after the review process to facilitate broader applications in robotics, such as simulation-driven learning or real-time control. Benchmark results demonstrate state-of-the-art performance, with timings ranging from 5 us for a 7-dof manipulator to 95 us for a 36-dof humanoid—an improvement of at least two orders of magnitude over alternative methods, such as automatic differentiation.
11footnotetext: Equal contribution.
I Introduction
I-A Context
Recent progress in reinforcement learning and trajectory optimization methods in robotics extensively relies on simulation. Additional information, such as the simulator derivatives of the simulator, might be leveraged to accelerate the convergence speed of these control methods. However, simulating robotics systems interacting with their environment induces a sequence of nonsmooth operations. Typically, collision detection involved in simulators is intrinsically nonsmooth (e.g., contact points might jump from one vertex to another one when slightly changing the orientation of the geometries), and frictional contact dynamics corresponds to nonsmooth problems (e.g., when a cube switches from a sticking mode to a sliding mode).
Several approaches have been envisaged to estimate simulator derivatives. Mordatch et al. [34] leverages MuJoCo [42] and finite differences to discover new behaviors. However, computing gradients via finite differences requires as many calls to forward dynamics as the number of parameters to differentiate, which becomes quickly prohibitive. Following the advent of the differentiable programming paradigm, DiffTaichi [25] and NeuralSim [23] propose to exploit Automatic Differentiation to differentiate through simplified contact models and geometries. In this vein, Brax [17] and MuJoCo MJX build on JAX [5] auto-diff and hardware acceleration capabilities to compute gradients through the computational graph. Because collision detection and contact forces involve iterative algorithms, the cost of computing gradients scales with the number of iterations performed during the evaluation of the forward dynamics. Alternatively, inspired by differentiable optimization [4], multiple works propose to apply implicit differentiation to a linear complementarity problem (LCP) [10, 18, 46] or mixed linear complementarity problem (MLCP) [38] relaxing the original nonlinear complementarity problem (NCP). Implicit differentiation has been extended to the NCP case in [24, 29], but it remains inefficient as it does not, for instance, exploit the structure induced by the kinematic chain. Other approaches, such as [47], rely on compliant contact models and focus on differentiating with respect to morphological parameters. For a comprehensive study of different contact models and the impact of relaxing the original NCP on gradients, we refer the reader to [30, 50].
I-B Contributions
In this paper, we present a comprehensive framework for differentiable simulation that combines differentiable rigid-body dynamics, differentiable collision detection, and differentiable contact resolution. We notably introduce an implicit differentiation scheme to compute the gradients of the NCP associated with the frictional contact problem without any relaxation and chain it with rigid body dynamics algorithms to finely exploit the kinematic sparsity of the problem. Our approach achieves substantial computational speedups, with gradient computations up to 100 times faster than current state-of-the-art methods in robotics, while avoiding any physical relaxation or geometrical approximation on meshes. We validate the effectiveness of our method by applying it to complex inverse problems, including the estimation of initial conditions and inverse dynamics through contact. We also show that our gradients can be used in a policy learning context to improve the sample efficiency during training. To support reproducibility and further research, we will make our code publicly available after the review process.
I-C Paper organization
In Sec. II, we provide a background on collision detection, frictional dynamics, rigid multi-body dynamics, and implicit differentiation techniques. Sec. III corresponds to the core contribution of this paper. We specify the computational graph of a physics engine and explain how the combination of gradients of rigid-body dynamics, collision detection, and contact forces make simulation end-to-end differentiable. Additionally, we show how implicit differentiation and rigid body algorithms can be leveraged to compute the derivatives of multibody frictional contact problems efficiently, including collision geometry contributions. In Sec. IV, the efficiency of our approach is benchmarked on several advanced robotics systems. We also leverage our differentiable physics engine to tackle various estimation and control problems and to efficiently learn policies. Sec. V discusses this work’s limitations and how it could set the stage for future developments in model-based approaches for robotics.
TABLE I: Differentiable physics engines for robotics.
[…]
II Background
This section reviews the three fundamental stages of modern robotic simulators: collision detection, contact modeling and multibody dynamics. We also review the notion of implicit differentiation, which is at the core of our approach.
II-A Collision detection
The collision detection phase identifies the contact points between the colliding geometries composing a simulation scene. Given two shapes and their relative poses, a collision detection algorithm (e.g., GJK [20] combined with EPA [44]) computes a contact point and a contact normal, corresponding to the direction separating the two bodies with minimal displacement. We define the contact frame $c$ with its origin at the contact point, and the Z axis aligned with the contact normal. Collision detection algorithms often assume the geometries to be convex but existing algorithms [45] can be employed during an offline preprocessing phase to decompose the nonconvex shapes into convex sub-shapes.
Collision detection is inherently nonsmooth for non-strictly convex geometries [15]. Concretely, this induces discontinuous contact points and normals. Thus, differentiating the contact point, the contact normal, and the contact frame w.r.t. the body poses is challenging. [43] uses a smooth approximation of the bodies to calculate the contact frame Jacobians. In contrast, [33] employs a randomized smoothing approach to compute the derivatives of contact points and normals.
II-B Frictional contact dynamics
Given a contact frame between two bodies, let $\bm{\lambda}$ denote the contact force and $\bm{\sigma}$ the contact velocity. The Signorini condition provides a complementarity constraint $0\leq\bm{\lambda}_{N}\perp\bm{\sigma}_{N}\geq 0$, ensuring the normal force is repulsive, bodies do not interpenetrate further, and no simultaneous separation motion and contact force exist. The maximum dissipation principle (MDP) combined with the frictional Coulomb law $\|\bm{\lambda}_{T}\|\leq\mu\lambda_{N}$ of friction $\mu$ states that $\bm{\lambda}_{T}\in\operatorname{argmax}_{y,\|y\|\leq\mu\bm{\lambda}_{N}}-y^{\top}\bm{\sigma}_{T}$ maximizes the power dissipated by the contact. These three principles are equivalent to the following nonlinear complementarity problem (NCP)
$\mathcal{K}_{\mu}$ $\ni\bm{\lambda}\perp\bm{\sigma}+\Gamma_{\mu}(\bm{\sigma})\in\mathcal{K}_{\mu}^{*},$ (1)
$\bm{\sigma}$ $=G\bm{\lambda}+g,$
where $G$ is the so-called Delassus matrix [12] that gives the system inverse inertia projected on the contacts. It is a linear operator mapping contact forces to contact velocities. $\bm{g}$ is the free velocity of the contact. $\mathcal{K}_{\mu}$ is a second-order cone with aperture angle $\operatorname{atan}(\mu)$, $\mathcal{K}_{\mu}^{*}=\mathcal{K}_{1/\mu}$ its dual cone and $\Gamma_{\mu}(\bm{\sigma})=[0,0,\mu\|\bm{\sigma}_{T}\|]$ is the so-called De Saxcé correction [1, 11] enforcing the Signorini condition [1, 30]. Problem (1) can be solved by interior point methods [24], projected Gauss-Seidel [27] or ADMM-based approaches [1, 8, 41].
II-C Multibody frictional contact dynamics
We briefly introduce the simulation of rigid bodies in contact, a core component of physics engines. We refer to [30] for a more detailed background. Let $\bm{q}\in\mathcal{Q}\cong\mathbb{R}^{n_{q}}$ denotes the joint position vector with $\mathcal{Q}$ the configuration space, i.e., the space of minimal coordinates. The equations of a constrained motion writes
$M(\bm{q})\bm{\dot{v}}+b(\bm{q},\bm{v})-\bm{\tau}=J_{c}^{\top}(\bm{q},c(\bm{q}))\bm{\lambda},$ (2)
where we denote by $\bm{v}\in\mathcal{T}_{q}\mathcal{Q}\cong\mathbb{R}^{n_{v}}$ and $\bm{\tau}\in\mathcal{T}_{q}^{*}\mathcal{Q}\cong\mathbb{R}^{n_{v}}$ the joint velocity vector and the joint torque vector. $\dot{\bm{v}}$ is the time derivative of $\bm{v}$. $M(\bm{q})$ is the joint-space inertia matrix, and $b(\bm{q},\bm{v})$ includes terms related to the gravity, Coriolis, and centrifugal effects. $J_{c}(\bm{q},c(\bm{q}))$ is the contact Jacobian associated with the contact frame $c(\bm{q})$ given by the collision detection on the system bodies using the configuration $\bm{q}$. In the following, we drop the dependency on the parameters when it is explicit.
To deal with rigid-body dynamics and impacts, we use an impulse-based formulation [32] obtained with the Euler symplectic scheme
$\bm{v}^{+}=\bm{v}+\Delta t\left(\bm{\dot{v}}_{f}+M^{-1}J_{c}^{\top}\bm{\lambda}\right),$ (3)
where $\bm{\dot{v}}_{f}=M^{-1}\left(\bm{\tau}-b\right)$ is the free acceleration term and $\Delta t$ is the time step. The acceleration term $\bm{\dot{v}}_{f}+M^{-1}J_{c}^{\top}\bm{\lambda}$ of (3) correspond to the unconstrained forward dynamic (UFD) with exterior forces $\bm{\lambda}$ which can be efficiently computed with the Articulated Body Algorithm (ABA) [16]. In the remaining of the paper, we use the shorthand $\text{UFD}(\bm{q},\bm{v},\bm{\tau},\bm{\lambda})=\bm{\dot{v}}_{f}+M^{-1}J_{c}^{\top}\bm{\lambda}$. Multiplying Eq. (3) by $J_{c}$, we recover the contact velocity associated to the contact NCP (1) in the case of multibody dynamics:
$J_{c}\bm{v}^{+}=\bm{\sigma}=G\bm{\lambda}+g.$ (4)
This yields the expression of the Delassus matrix $G=J_{c}M^{-1}J_{c}^{\top}$ and the free contact velocity vector $g=J_{c}(\bm{v}+\Delta t\bm{\dot{v}}_{f})$. For poly-articulated rigid-body systems, $G$ depends on $\bm{q}$, and $g$ depends on $\bm{q},\bm{v},\bm{\tau}$. In this respect, the associated NCP is conditioned by $\bm{q},\bm{v},\bm{\tau}$. Note also that the contact Jacobian $J_{c}$ depends on $\bm{q}$, first through the kinematic structure of the system and second through the contact frame $c(\bm{q})$.
II-D Implicit differentiation
As previously mentioned, the physically accurate contact forces denoted by $\bm{\lambda}^{*}$ are implicitly defined as the solution of the NCP (1), we write $0=\text{NCP}(\bm{\lambda}^{*};\bm{q},\bm{v},\bm{\tau})$. By deriving the optimality conditions, the implicit function theorem allows the computation of their gradients and corresponds to the concept of implicit differentiation [4]. The theorem provides locally the derivatives of the solution $\text{d}\bm{\lambda}^{*}$ as a linear function of the other variable derivatives.
This approach has been successfully applied to differentiate Quadratic Programming (QP) problems in [3] and generalized to convex cone programs [2] and LCPs [10]. More generally, it allows incorporating optimization layers in the differentiable programming paradigm. In Sec. III-B, we extend this approach to the NCP case and propose a method to compute the gradients of the contact forces efficiently.
III Efficient differentiable simulation
This section details the core contribution of this paper, namely a comprehensive framework for differentiable simulation that combines differentiable rigid-body dynamics, differentiable collision detection, and differentiable contact resolution. We show the link between the derivatives of multibody dynamics, frictional contact dynamics, and collision detection. We introduce an efficient algorithmic solution to compute the derivatives associated with the contact NCP by solving a reduced system of equations of minimal dimension resulting from its implicit differentiation.
III-A Chaining rigid-body dynamics derivatives and NCP derivatives
From Sec. II, the simulation equations (3) can be restated using unconstrained forward dynamic (UFD) and the solution to the (NCP) problem (1) as
00 0 $=\text{NCP}(\bm{\lambda}^{*};\bm{q},\bm{v},\bm{\tau})$ (5)
$\bm{v}^{+}$ $=\bm{v}+\text{UFD}(\bm{q},\bm{v},\bm{\tau},\bm{\lambda}^{*})\Delta t.$ (6)
Next, we consider forward-mode differentiation setting [21], i.e., we aim to compute the Jacobian $\frac{\text{d}v^{+}}{\text{d}\theta}$ where $\theta\in\mathbb{R}^{n_{p}}$ represents any subset of the inputs $\{\bm{q},\bm{v},\bm{\tau}\}$ or physical parameters. Our approach can be efficiently adapted to the reverse mode by applying the computational trick introduced in [3]. Differentiating (6) leads to
$\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}=$ $\frac{\partial\text{UFD}}{\partial\bm{\lambda}}\frac{\text{d}\bm{\lambda}^{*}}{\text{d}\theta}\Delta t$ (7)
$+\underbrace{\frac{\text{d}\bm{v}}{\text{d}\theta}+\left(\frac{\partial\text{UFD}}{\partial\bm{q}}\frac{\text{d}\bm{q}}{\text{d}\theta}+\frac{\partial\text{UFD}}{\partial\bm{v}}\frac{\text{d}\bm{v}}{\text{d}\theta}+\frac{\partial\text{UFD}}{\partial\bm{\tau}}\frac{\text{d}\bm{\tau}}{\text{d}\theta}\right)\Delta t}_{\left.\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}\right|_{\lambda=\lambda^{*}}},$
where we identify the term $\left.\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}\right|_{\lambda=\lambda^{*}}$ of derivatives, considering that $\bm{\lambda}^{*}$ does not vary. The derivatives $\frac{\partial\text{UFD}}{\partial\bm{q},\bm{v},\bm{\tau}}$ and $\frac{\partial\text{UFD}}{\partial\bm{\lambda}}=M^{-1}(\bm{q})J_{c}^{\top}(\bm{q})$ can be efficiently computed via rigid-body algorithms [6] and are, for instance, available in Pinocchio [7]. At this stage, it is worth noting that $\frac{\partial\text{UFD}}{\partial\bm{q}}$ also depends on the geometry of the contact through the contact Jacobian $J_{c}(\bm{q},c(\bm{q}))$. The classical ABA rigid body algorithm gives the derivatives related to the first variable, and we need to add a term related to the variations of $J_{c}$ induced by the variation of the contact point $c(\bm{q})$ (see Section III-D and Appendix C).
Computing the sensitivity of the contact forces $\frac{\text{d}\bm{\lambda}^{*}}{\text{d}\theta}$ is also challenging as $\bm{\lambda}^{*}$ is obtained implicitly by solving a NCP (1) which depends on $\bm{q}$, $\bm{v}$ and $\bm{\tau}$ through $G$ and $g$. Notably, the solutions of the NCP are intrinsically nonsmooth, corresponding to the solution of a differential inclusion problem [1]. To understand the nonsmoothness of the NCP solutions, consider the case of a contact force either (i) saturating the Coulomb cone or (ii) lying strictly inside the cone. In case (i), the contact force variations must lie on the tangent plane to the cone at this force value, while in case (ii), no restriction on the contact force variations applies. The derivatives do not lie on the constraint manifolds in these two cases. Next, we detail how implicit differentiation of (5) can be leveraged to compute them precisely at a limited computational cost.
III-B Implicit differentiation of the NCP
The dynamics induced by the NCP (1) is inherently nonsmooth as it can switch on three modes. These modes correspond to the active set of (1) and result in different gradients for the contact dynamics. Our approach considers scenarios with multiple contact points and requires the identification of the mode for each contact. For clarity purposes, we present the equations for a single contact point in the case of each of the three modes.
Mode 1 - Breaking contact (brk). This mode corresponds to the case where the contact is separating ($\bm{\sigma}_{N}>0$), which is induced by the Signorini condition that $\bm{\lambda}^{*}=0$. This mode can be treated separately from the other two modes, as the contact force is zero and the contact point velocity is not constrained, yielding
$\frac{\mathrm{d}\bm{\lambda}^{*}}{\text{d}\theta}=0.$ (8)
Mode 2 - Sticking contact (stk). In this mode, the contact point is not moving ($\bm{\sigma}=0$), which yields the same equations as a bilateral constraint of an attached point $G\bm{\lambda}^{*}+g=0$. Differentiating this constraint gives the following linear equations on $\frac{\mathrm{d}\bm{\lambda}^{*}}{\text{d}\theta}$
$G\frac{\mathrm{d}\bm{\lambda}^{*}}{\text{d}\theta}=-\left(\frac{\text{d}G}{\text{d}\theta}\bm{\lambda}^{*}+\frac{\mathrm{d}g}{\text{d}\theta}\right).$ (9)
Figure 1: Illustration of the sliding mode. $\bm{\lambda}^{*}$ lives in the boundary of the cone $\mathcal{K}_{\mu}$ in the direction opposite to $\bm{\sigma}=\bm{\sigma}_{T}$ and the variation $\text{d}\bm{\lambda}^{*}$ lies inside the tangent plane.
Mode 3 - Sliding contact (sld). In this regime, the contact point moves on the contact surface, which implies a null normal velocity ($\bm{\sigma}_{N}=0$) and a non-null tangential velocity ($\|\bm{\sigma}_{T}\|>0$). Moreover, from the MDP, tangential contact forces should lie on the boundary of the cone and in the opposite direction of the tangential velocity ($\bm{\lambda}^{*}_{T}=-\mu\bm{\lambda}^{*}_{N}\frac{\bm{\sigma}_{T}}{\|\bm{\sigma}_{T}\|}$). This additionally implies that $\frac{\mathrm{d}\bm{\lambda}^{*}}{\text{d}\theta}$ should be in the plane tangent to the friction cone as illustrated in Fig. 1 and allows reducing the search space to a 2D plane via a simple change of variable $\frac{\mathrm{d}\bm{\lambda}^{*}}{\text{d}\theta}=R\frac{\mathrm{d}\bm{\tilde{\lambda}}}{\text{d}\theta}$ where $R=\begin{pmatrix}\frac{\bm{\lambda}}{\|\bm{\lambda}\|}&\bm{e}_{z}\times\frac{\bm{\sigma}_{T}}{\|\bm{\sigma}_{T}\|}\end{pmatrix}\in\mathbb{R}^{3\times 2}$. Therefore, differentiating these equations yields the following conditions on the gradients
$\tilde{G}\frac{\mathrm{d}\bm{\tilde{\lambda}}}{\text{d}\theta}=-{R}^{\top}P\left(\frac{\mathrm{d}G}{\text{d}\theta}\bm{\lambda}^{*}+\frac{\mathrm{d}g}{\text{d}\theta}\right),$ (10)
where: $P=\begin{pmatrix}H(\bm{\sigma}_{T})&0_{2\times 1}\\ 0_{1\times 2}&1\end{pmatrix}\in\mathbb{R}^{3\times 3}$, $H(x)=\frac{1}{\alpha}\left(\text{Id}-\frac{x}{\|x\|}\frac{x}{\|x\|}^{\top}\right)\in\mathbb{R}^{2\times 2}$, $\tilde{G}={R}^{\top}PG{R}+Q$ with $Q=\begin{pmatrix}0&0\\ 0&1\end{pmatrix}\in\mathbb{R}^{2\times 2}$ and $\alpha=\frac{\|\bm{\sigma}_{T}\|}{\mu\bm{\lambda}_{N}}$. We refer to A for the detailed derivation.
III-C Efficient computation: exploiting kinematic-induced sparsity
First, one should identify the active contact modes to obtain the equations for all contact points. We denote $\mathcal{A}_{\text{brk}}$, $\mathcal{A}_{\text{stk}}$, and $\mathcal{A}_{\text{sld}}$ as the sets of contact indices corresponding to the breaking, sticking, and sliding contacts respectively. The dynamics of the different contacts are coupled through the Delassus matrix $G$. Thus, we construct a matrix $A\in\mathbb{R}^{(3n_{stk}+2n_{sld})\times(3n_{stk}+2n_{sld})}$ where from $G$ we remove the blocks related to $\mathcal{A}_{\text{brk}}$ and modify the lines and columns of $G$ related to $\mathcal{A}_{\text{sld}}$ by following the pattern of $\tilde{G}$ presented in (10). Once this is done, the reduced linear system on $X$, the stacking of $\text{d}\bm{\lambda}^{*}$ and $\text{d}\tilde{\bm{\lambda}}$, is obtained by concatenating the corresponding right-hand side of (9) and (10). We obtain the linear system corresponding to implicit differentiation
$AX=-B\left(\frac{\mathrm{d}G}{\text{d}\theta}\bm{\lambda}^{*}+\frac{\mathrm{d}g}{\text{d}\theta}\right),$ (11)
where $B$ is block diagonal with identity for blocks of $\mathcal{A}_{\text{stk}}$ and the basis change $R^{\top}P$ for blocks of $\mathcal{A}_{\text{sld}}$; the complete construction is given in the Appendix B.
Computing the right-hand side of (11) requires evaluating the derivative of $G\bm{\lambda}^{*}+g$ with $\bm{\lambda}^{*}$ taken constant:
$\frac{\mathrm{d}G}{\text{d}\theta}\bm{\lambda}^{*}+\frac{\mathrm{d}g}{\text{d}\theta}=\left.\frac{\text{d}G\bm{\lambda}^{*}+g}{\text{d}\theta}\right|_{\lambda=\lambda^{*}}.$ (12)
By recalling that $G\bm{\lambda}^{*}+g=J_{c}\bm{v}^{+}$ (see Eq. 56)is the contact point velocity, this term exactly corresponds to the derivatives w.r.t. $\theta$ of the contact point velocity with $\bm{\lambda}^{*}$ taken constant. Calculating the derivative, we obtain:
$\frac{\text{d}G}{\text{d}\theta}\bm{\lambda}^{*}+\frac{\mathrm{d}g}{\text{d}\theta}=J_{c}\left.\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}\right|_{\lambda=\lambda^{*}}+\left.\frac{\text{d}J_{c}\bm{v}^{+}}{\text{d}\theta}\right|_{v=v^{+}}.$ (13)
The first term is already computed thanks to the ABA derivatives (7[6], and the second term $\left.\frac{\text{d}J_{c}\bm{v}^{+}}{\text{d}\theta}\right|_{v=v^{+}}$, which is the derivatives of the contact velocity with $\bm{v}^{+}$ assumed to be constant, can be computed at a reduced cost via the partial derivatives of the forward kinematics [6] evaluated in $\bm{q},\bm{v}^{+}$. This allows us to avoid the expensive computation of $\frac{\mathrm{d}G}{\text{d}\theta}$ as it is a tensor in general, while only its product with $\bm{\lambda}$ is required. It is worth noting at this stage that the term $\left.\frac{\text{d}J_{c}\bm{v}^{+}}{\text{d}\theta}\right|_{v=v^{+}}$ also depends on the geometry of the contact. Therefore, computing gradients w.r.t. to the configuration $\bm{q}$ requires evaluating an additional term for the variations of $J_{c}$ induced by the variations of contact points on the local geometries [33] and presented in the next subsection.
Finally, to obtain $\frac{\text{d}\bm{\lambda}^{*}}{\text{d}\theta}$, we solve the linear system (11) using a QR decomposition of $A$ before projecting back the reduced variables $\text{d}\tilde{\bm{\lambda}}$ in $\mathbb{R}^{3}$. At this stage, it is worth noting that these gradients are computed given the current active set, and thus they do not capture the information on the contact modes boundary. Still, this is possible by combining our approach with smoothing techniques explored in [39, 36, 49].
III-D Collision detection contribution
The collision detection phase depends on the body poses induced by the configuration $\bm{q}$. Thus, when $\theta$ depends on $\bm{q}$, one must consider the variation of the contact location given variations of $\bm{q}$. Recent work on differentiable collision detection [33, 43] allows computing the derivative of the normal and contact point w.r.t. the poses of the bodies.
By choosing a function that constructs a contact frame $c$ from a contact point and its normal, we compute the derivative of this frame w.r.t. the body poses. Chaining this derivative with the usual kinematics Jacobian, which relates the variation of $\bm{q}$ to the variation of body poses, one can obtain $\frac{\text{d}c}{\text{d}\theta}$. The frame $c$ intervenes in $J_{c}$ through a change of frame (the adjoint of the placement). By leveraging spatial algebra[16] (see Appendix C for the details), we calculate $\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}$ and $\frac{\partial J_{c}\bm{v}^{+}}{\partial c}$. We add the first collision term $\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ in $\left.\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}\right|_{\lambda=\lambda^{*}}$ and the second collision term $\frac{\partial J_{c}\bm{v}^{+}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ in $\left.\frac{\text{d}J_{c}\bm{v}^{+}}{\text{d}\theta}\right|_{v=v^{+}}$ to account for the variation of $J_{c}$ due to the variation of the contact points.
The complete details of the terms and simulation derivative $\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}$ are reported in Appendix D.
IV Experiments
In this section, we first demonstrate state-of-the-art computational timings when computing simulation gradients for various robotic systems (Fig. 2). Second, we apply our approach to solve two inverse problems involving contact dynamics: retrieving an initial condition that leads to a target final state and finding a torque that yields a null acceleration on a quadruped. Eventually, we use our approach in first-order policy learning algorithms in order to efficiently train control policies on various systems. The experiments as well as additional experiments presented in Appendix E are available in the video attached.
Figure 2: The robotics systems used to evaluate our approach range from simple systems such as MuJoCo’s half-cheetah (Left) to more complex high-dof robots such as Unitree’s Go1 (Center) and H1 (Right)
Implementation details. We have implemented our analytical derivatives in C++ for efficiency. We leverage open-source software of the community: Eigen [22] for efficient linear algebra, Pinocchio [7] for fast rigid body dynamics and their derivatives, and HPP-FCL [35] for high-speed collision detection. The code associated with this paper will be released as open-source upon acceptance. All the experiments are performed on a single core of an Apple M3 CPU.
IV-A Timings
[…]
TABLE II: Details of scenarios for computational timings. For the number of vertices per mesh, N/A indicates a scenario that does not contain any mesh.
[…]
TABLE III: Comparative analysis between ours, MuJoCo, and MJX frameworks. Timing statistics (mean and standard deviation in microseconds) for simulation, gradient, and finite-differences computation for one simulation step. For MJX, N/A denotes scenarios where geometries were not supported.
The computational efficiency of our approach is evaluated by measuring the average time required to compute the full Jacobian of the simulator, i.e., $\frac{\text{d}\bm{v}^{+}}{\text{d}\bm{q},\bm{v},\bm{\tau}}$, along a trajectory. We consider various scenarios ranging from simple systems composed of basic geometry primitives (MuJoCo’s half-cheetah and humanoid) to more complex and realistic robots with multiple DoFs and complex geometries (UR5, Unitree Go1, and Boston Dynamics Atlas). Tab. II reports the numbers associated with the different robots considered. To stabilize the simulation behaviors, we compute contact collision patches (composed of 4 contact points each), thus substantially increasing the dimensions of the problem to solve.
Tab. III demonstrates computational timings for gradient computation that are of the same order of magnitude as simulation and significantly faster than central finite differences. As another point of comparison, Nimble [46] requires 1ms and 16ms on half-cheetah and Atlas, corresponding to an approximate speedup factor of 100 for our method. In Tab. III, we also compare our approach to MuJoCo MJX GPU simulation pipeline. The numbers for gradients computed on GPU correspond to the samples generated after one second when using all the threads of a Nvidia A100. We find our approach to be competitive even though it operates on a single CPU core. Importantly, our performance gain is obtained although we work on the unrelaxed NCP and with full meshes descriptions (cf. Tab. I). In general, this is not achieved by current GPU simulation approaches which have to operate on relaxed physical principles and simplified geometries due to hardware constraints.
IV-B Inverse problems
Figure 3: Estimation of initial conditions. A Gauss-Newton (GN) algorithm can leverage the efficient implicit differentiation to accurately retrieve the initial velocity $\bm{v}_{0}$ and impulse $\bm{\tau}_{0}$. On the third and fourth figures, the black curve representing Gradient Descent with finite differences rises due to the excessively large estimated gradients. When at the boundary of a contact mode, the norm of the finite differences gradients becomes inversely proportional to the step size used.
Estimating initial conditions. As a first application of differentiable simulation, we aim at retrieving the initial condition $\theta$ (either the initial velocity $\bm{v}_{0}$ or an initial impulse $\bm{\tau}_{0}$), leading to a target final state $\bm{q}_{T}^{*}$ after $T$ time steps. Here, we consider the case of a cube thrown on the floor evolving in a sliding mode. The problem can be written as:
$\min_{\bm{\theta}}\frac{1}{2}\left\|\bm{q}_{T}(\theta)-\bm{q}_{T}^{*}\right\|_{2}^{2},$ (14)
where $\bm{q}_{T}$ is the final configuration and depends on the initial velocity and impulse. Our forward-mode differentiation allows us to efficiently compute the Jacobian of $\bm{q}_{T}$ w.r.t. $\bm{\theta}$. We leverage this feature to implement a Gauss-Newton (GN) approximation of the Hessian of (14). Fig. 3 demonstrates the benefits of using our implicit gradients over finite-differences in order to reach a precise solution of (14). Moreover, exploiting the full Jacobian in a quasi-Newton algorithm also reduces the number of iterations compared to a vanilla gradient descent.
Retrieving the initial impulse on the cube $\bm{\tau}_{0}$ is a challenging nonsmooth and nonconvex optimization problem, which can explain the plateau reached by our vanilla Gauss-Newton implementation. Working on dedicated nonsmooth optimization algorithms is a promising research direction that could lead to higher-quality solutions.
Inverse dynamics through contacts. We evaluate our approach on an Inverse Dynamics (ID) task involving contacts. In particular, we aim at finding the torque on actuators $\bm{\tau}_{\text{act}}$ leading to a null acceleration for a Unitree Go1 quadrupedal robot in a standing position $(\bm{q},\bm{v})$. By denoting $S$ the actuation matrix, the ID problem can be formulated as follows:
$\min_{\bm{\tau}_{\text{act}}}\frac{1}{2}\left\|\bm{v}^{+}(\bm{q},\bm{v},S^{\top}\bm{\tau}_{\text{act}})-\bm{v}^{*}\right\|_{2}^{2},$ (15)
where the initial $\bm{v}$ and target $\bm{v}^{*}$ velocities are null in this example. As previously explained, we use the Jacobians computed by our differentiable simulator with a Gauss-Newton algorithm. As shown by Fig. 4, the problem is solved with high accuracy in only a few iterations (approx. 10 to reach an error of $51\mathrm{e}-$ 5). Just like in the initial conditions estimation setup, implicit gradients allow us to solve the problem with a higher precision than finite differences.
Figure 4: Contact inverse dynamics on an underactuated Go1 quadruped can be efficiently performed via a Gauss-Newton algorithm by leveraging the derivatives of our differentiable simulator.
IV-C Applications to Policy Learning
Model-free reinforcement learning relies on zeroth-order gradient estimation via the policy gradient theorem, often leading to high variance and slow convergence. In contrast, first-order gradient-based optimization using analytical gradients from a differentiable simulator enables more efficient policy updates e.g. SHAC [48] and AHAC [19].
Figure 5: SHAC vs. PPO on cartpole swing up (upper) and hopper (lower). SHAC algorithm leverages differentiable simulation to achieve improved sample efficiency.
We integrate our differentiable simulator with the first-order, on-policy algorithm SHAC and compare it to the zeroth-order, on-policy algorithm PPO. We evaluate two illustrative systems: the cartpole swing-up and the MuJoCo hopper. The performances are reported in Fig.  5. While these systems are relatively small, they are not trivial due to their non-smooth dynamics. The cartpole involves joint limits, and the hopper combines joint constraints with ground contact interactions, making these tasks representative of the complexities encountered in real-world scenarios. Commonly used robotics benchmarks in RL are from Gymnasium and rely on the MuJoCo simulator [42], which relaxes contact constraints using compliance [30]. First-order RL benchmarks with SHAC or AHAC work with even more relaxed dynamics by using the dflex simulator. The later models all constraints with a spring-damper approximation which results in smoother dynamics. In contrast, our simulation approach allows modeling non-compliant contacts and joint limits without any such relaxations [8]. In turn, the absence of relaxation directly affects the smoothness and continuity of the gradients, often making it challenging to leverage them for gradient based optimization. Certain reward terms in Gymnasium environments are non-smooth and require smoothing, as proposed in [48], to allow informative gradient computation (see section F in the Appendix).
The experiments show that using our differentiable simulator with SHAC leads to greater sample efficiency compared to PPO. However, the training process is noticeably noisier. This may arise from the stiffness of our simulator, which models non-smooth constraints directly rather than using the commonly employed relaxations or spring-damper systems. Although differentiable simulation reduces gradient variance, the resulting gradients can become highly unstable in stiff scenarios. Mitigating this instability is a challenging problem [19] and may require effective gradient smoothing strategies. Potential approaches include adaptive scheduling of a compliance term for the constraints or leveraging solver methods that inherently apply smoothing, such as interior point solvers [37, 24], to enhance stability without sacrificing efficiency.
V Limitations
This paper introduces an end-to-end differentiable physics pipeline for robotics based on the implicit differentiation of the non-relaxed NCP for contacts. By avoiding any relaxation, we prevent the appearance of nonphysical simulation artifacts. Moreover, exploiting the sparsity induced by the robot kinematic chains and leveraging the derivatives of rigid body algorithms allows us to achieve state-of-the-art timings, with a speed-up of at least $100$ compared to alternative solutions of the state of the art. In an MPC context where the dynamics and its derivatives are evaluated at a high frequency, the gains in physical realism and efficiency could determine the controller’s overall performance.
Yet, as the NCP induces inherently non-smooth dynamics, exploiting its gradients requires dedicated algorithms when addressing downstream optimization tasks. Our experiments with first-order reinforcement learning highlight this challenge: directly using the simulator’s raw analytical gradients with NCP-modeled constraints results in efficient but sometimes less stable convergence. Some previous works [40, 31, 36] leverage randomized smoothing techniques that provide smooth gradient estimates from simulation and gradient samples. Alternative solutions relax the physics, either explicitly [42, 34, 28] or implicitly, by leveraging interior-points (IP) methods [24].
Until now, none of the previous first-order RL works [48, 19] showed a sim-to-real transfer on real robotic hardware. This might be due to the significant difference between the smoothed simulators and reality. Therefore, working with the challenging-to-use but more accurate non-smooth gradients from more accurate simulators might pave the way to a sim-to-real transfer.
Similarly to existing robotic simulators (e.g., MuJoCo, Bullet, DART), this paper models contact interactions as vanilla 3D contact points, while richer but more complex contact models exist. One promising research direction could consider extending this work towards deformable contact interactions to enhance simulator realism, such as in [14].
VI Conclusion
In future work, we plan to combine our qualitative gradient approach with smoothing techniques to ease the integration within recent control frameworks to tackle more complex robotics tasks. Specifically, this paper introduces an efficient method for computing physics gradients, which are used in first-order policy learning algorithms such as SHAC [48] or AHAC[19]. The next step is extending our SHAC experiments to more complex real-world robotics systems. It is also worth noticing that our approach is not limited to rigid-body robots, but could also be leveraged for soft dynamics in general to design and control soft robots [13] and could be adapted to use implicit integrator as in [9]. The proposed differentiable simulation of an unrelaxed physics model is a crucial step toward reducing the Sim2Real gap [26], and future work may adapt learning algorithms to effectively achieve this goal. Finally, we hope this work will serve as a catalyzer in the robotics and learning communities and motivate the development of new reinforcement learning and trajectory optimization methods leveraging simulation gradients in order to accelerate the discovery of complex robot movements in contact.
Appendix A Sliding mode
In this section, we detail the equations (Eq. 10 of the paper) of the sensitivity analysis of the contact force in the case of a sliding mode ($\|\bm{\sigma}_{T}\|>0$). In sliding mode, both the contact forces and the contact point velocity are on the border of their cone. Thus, from the NCP we have:
$\|\bm{\lambda}_{T}\|=\mu\bm{\lambda}_{N}$ (16a)
$\bm{\sigma}_{N}=0$ (16b)
${\bm{\lambda}}^{\top}\big{(}\bm{\sigma}+\Gamma_{\mu}(\bm{\sigma})\big{)}=0$ (16c)
$\bm{\sigma}=G\bm{\lambda}+g,$ (16d)
which is equivalent to:
$\bm{\lambda}_{T}=-\mu\bm{\lambda}_{N}\frac{\bm{\sigma}_{T}}{\|\bm{\sigma}_{T}\|}$ (17a)
$\bm{\sigma}_{T}=G_{T}\bm{\lambda}+g_{T}$ (17b)
$G_{N}\bm{\lambda}+g_{N}=0.$ (17c)
By differentiating, we get the following system on the derivatives:
$K\mathrm{d}\bm{\lambda}=-\frac{1}{\alpha}Hd\bm{\sigma}_{T}$ (18a)
$d\bm{\sigma}_{T}=G_{T}d\bm{\lambda}+(dG_{T}\bm{\lambda}+dg_{T})$ (18b)
$G_{N}\mathrm{d}\bm{\lambda}=-\left(\mathrm{d}G_{N}\bm{\lambda}+\mathrm{d}g_{N}\right),$ (18c)
where $K=\begin{pmatrix}\text{Id}_{2}&\mu\bm{u}_{T}\end{pmatrix}\in\mathbb{R}^{2\times 3}$ and $H=\left(\text{Id}-\bm{u}_{T}\bm{u}_{T}^{\top}\right)\in\mathbb{R}^{2\times 2}$ with $\bm{u}_{T}=\frac{\bm{\sigma}_{T}}{\|\bm{\sigma}_{T}\|}$ and $\alpha=\frac{\|\bm{\sigma}_{T}\|}{\mu\bm{\lambda}_{N}}$. We rewrite as
$\left(\frac{1}{\alpha}HG_{T}+K\right)\mathrm{d}\bm{\lambda}=-\frac{1}{\alpha}H\left(\mathrm{d}G_{T}\bm{\lambda}+\mathrm{d}g_{T}\right)$ (19a)
$G_{N}\mathrm{d}\bm{\lambda}=-\left(\mathrm{d}G_{N}\bm{\lambda}+\mathrm{d}g_{N}\right).$ (19b)
Stacking the two equations yields:
$(PG+\tilde{K})\mathrm{d}\bm{\lambda}=-P(\mathrm{d}G\bm{\lambda}+\mathrm{d}g),$ (20)
where we introduce $P=\begin{pmatrix}\frac{1}{\alpha}H&0_{2\times 1}\\ 0_{1\times 2}&1\end{pmatrix}\in\mathbb{R}^{3\times 3}$ and $\tilde{K}=\begin{pmatrix}K\\ 0_{1\times 3}\end{pmatrix}\in\mathbb{R}^{3\times 3}$.
Because $\bm{\lambda}$ is constrained to stay on the boundary of the cone, its variations $\text{d}\bm{\lambda}$ live in the tangent 2D plane whose basis is $R=\begin{pmatrix}\frac{\bm{\lambda}}{\|\bm{\lambda}\|}&\bm{e}_{z}\times\frac{\bm{\sigma}_{T}}{\|\bm{\sigma}_{T}\|}\end{pmatrix}\in\mathbb{R}^{3\times 2}$. Applying the change of variable $\mathrm{d}\bm{\lambda}=R\mathrm{d}\tilde{\bm{\lambda}}$ and multiplying the previous system 20 by $R^{\top}$ allows getting a linear system of reduced dimension:
$\tilde{G}\text{d}\tilde{\bm{\lambda}}=-{R}^{\top}P\left(\mathrm{d}G\bm{\lambda}+\mathrm{d}g\right),$ (21)
where $\tilde{G}=R^{\top}PGR+Q\in\mathbb{R}^{2\times 2}$ with $Q=R^{\top}\tilde{K}R=\begin{pmatrix}0&0\\ 0&1\end{pmatrix}\in\mathbb{R}^{2\times 2}$. Hence the final expression when taking derivative w.r.t $\theta$ in the optimal force $\bm{\lambda}^{*}$
$\tilde{G}\frac{\mathrm{d}\bm{\tilde{\lambda}}}{\text{d}\theta}=-{R}^{\top}P\left(\frac{\mathrm{d}G}{\text{d}\theta}\bm{\lambda}^{*}+\frac{\mathrm{d}g}{\text{d}\theta}\right).$ (22)
Appendix B Implicit NCP gradient system
In this section, we detail the final system solved to compute the gradients of the NCP (equation (10) of the paper). Following the paper notations, we denote $\mathcal{A}_{\text{brk}}$, $\mathcal{A}_{\text{stk}}$, and $\mathcal{A}_{\text{sld}}$ as the sets of contact indices corresponding to the breaking, sticking, and sliding contacts respectively and $n_{\text{brk}}$, $n_{\text{stk}}$, and $n_{\text{sld}}$ their cardinals and $n=n_{\text{brk}}+n_{\text{stk}}+n_{\text{sld}}$ the total number of contacts. For the implicit gradients system, we removed the $\text{d}\bm{\lambda}$ variables associated with contacts in $\mathcal{A}_{brk}$ and sorted the remaining ones by putting first the components associated with the contacts in $\mathcal{A}_{stk}$ before the reduced ones $\text{d}\tilde{\bm{\lambda}}$ of $\mathcal{A}_{sld}$. Then the total variation of $\bm{\lambda}$ is $\text{d}\bm{\lambda}=CX\in\mathbb{R}^{3n}$ with
$X=\begin{pmatrix}\text{d}\bm{\lambda}^{(1)}\\ \vdots\\ \text{d}\bm{\lambda}^{(n_{\text{stk}})}\\ \text{d}\tilde{\bm{\lambda}}^{(1)}\\ \vdots\\ \text{d}\tilde{\bm{\lambda}}^{(n_{\text{sld}})}\end{pmatrix}\in\mathbb{R}^{3n_{stk}+2n_{sld}}$ (23)
$C=\left(\begin{array}[]{c|c}0_{3n_{\text{brk}},3n_{\text{stk}}}&0_{3n_{\text{brk}},2n_{\text{sld}}}\\ \hline\cr\text{Id}_{3n_{\text{stk}},3n_{\text{stk}}}&0_{3n_{\text{stk}},2n_{\text{sld}}}\\ \hline\cr 0_{3n_{\text{sld}},3n_{\text{stk}}}&\begin{array}[]{ccc}R^{(1)}&&\\ &\ddots&\\ &&R^{(n_{\text{sld}})}\end{array}\end{array}\right)\in\mathbb{R}^{3n\times(3n_{stk}+2n_{sld})}$ (30)
Note that in the sticking case, the right-hand side is $\text{d}G\bm{\lambda}+\text{d}g$ and the left-hand side is $G\text{d}\lambda$ and for the sliding mode, the right-hand side is multiplied by $R^{\top}P$ and the right-hand side is composed by $R^{\top}P$ and $R$ plus an additional term due to $Q$. So if we introduce $B$ and $A$ as
$B$ $=\left(\begin{array}[]{c|c|c}0_{3n_{\text{stk}},3n_{\text{brk}}}&\text{Id}_{3n_{stk},3n_{stk}}&0_{3n_{\text{stk}},3n_{\text{sld}}}\\ \hline\cr 0_{2n_{\text{sld}},3n_{\text{brk}}}&0_{2n_{\text{sld}},3n_{stk}}&\begin{array}[]{ccc}R^{(1)\top}P^{(1)}&&\\ &\ddots&\\ &&R^{(n_{\text{sld}})\top}P^{(n_{\text{sld}})}\end{array}\end{array}\right)\in\mathbb{R}^{(3n_{stk}+2n_{sld})\times 3n}$ (36)
$A$ $=BGC+\left(\begin{array}[]{c|c}0_{3n_{\text{stk}},3n_{\text{stk}}}&0_{3n_{\text{brk}},2n_{\text{sld}}}\\ \hline\cr 0_{2n_{\text{sld}},3n_{\text{brk}}}&\begin{array}[]{ccc}Q&&\\ &\ddots&\\ &&Q\end{array}\end{array}\right)\in\mathbb{R}^{(3n_{stk}+2n_{sld})\times(3n_{stk}+2n_{sld})}$ (42)
where $G$ is the complete Delassus matrix that induces coupling between the different contacts. We recover the linear system of the implicit gradient
$AX=-B(\text{d}G\lambda+\text{d}g),$ (43)
and, finally, the derivative w.r.t $\theta$ of the force taken in the optimal force $\lambda^{*}$ as
$\frac{\mathrm{d}\bm{\lambda}^{*}}{\text{d}\theta}=-CA^{-1}B\left(\frac{\text{d}G}{\text{d}\theta}\bm{\lambda}^{*}+\frac{\mathrm{d}g}{\text{d}\theta}\right).$ (44)
Details on implementation. In practice, the matrix $B$ and $C$ are not computed, but we directly work on $G$ and $(\text{d}G\lambda+\text{d}g)$ by discarding the right lines and modifying groups of columns and lines to exploit the sparsity of $B$ and $C$. We compute the matrix $A$ and its inverse using a QR decomposition.
Appendix C Collision detection contribution
Here, we present the terms $\frac{\partial J_{c}\bm{v}^{+}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ and $\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ from Section III-D. Given a contact frame $c$ between body $1$ and body $2$, the contact Jacobian is
$J_{c}$ $=E({}^{c}X_{1}J_{1}-{}^{c}X_{2}J_{2}),$ (45)
where $E=\begin{pmatrix}\text{Id}_{3}&0_{3,3}\\ 0_{3,3}&0_{3,3}\end{pmatrix}$ is an operator that allows the extraction of the linear part, and $J_{1}$, $J_{2}$ are the kinematic Jacobian of the bodies $1$ and $2$. In practice, the contact placement $c$ is calculated relative the world frame using HPP-FCL. The contact placement is given in function of the placement of two bodies placement relative to the world: ${}^{0}M_{1}(\bm{q})$ and ${}^{0}M_{2}(\bm{q})$. We can write:
${}^{0}M_{c}(\bm{q})$ $={}^{0}\text{CD}({}^{0}M_{1}(\bm{q}),{}^{0}M_{2}(\bm{q})),$ (46)
where ${}^{0}\text{CD}$ is an acronym for collision detection. We are interested in the derivatives of $J_{c}\bm{v}$ when $J_{1}$ and $J_{2}$ are considered constant because their derivation is already considered in the other terms. For clarity of the presentation, we present the calculation for ${}^{c}X_{1}J_{1}v$ as the derivation for $J_{c}v$ follows naturally. In term of Lie groups, ${}^{c}X_{1}=\operatorname{Ad}_{{}^{0}M_{c}^{-1}(\bm{q}){}^{0}M_{1}(\bm{q})}$ where $\operatorname{Ad}$ denotes the adjoint operator on $\mathrm{SE}(3)$ to explicitly show the dependency in the variables. With the rules of spatial algebra, we have for a vector $x$ of the lie algebra, and placement $M$, $M^{\prime}$ in $\mathrm{SE}(3)$
$\text{d}(\operatorname{Ad}_{M}x)$ $=-\operatorname{ad}_{\operatorname{Ad}_{M}x}\operatorname{Ad}_{M}\text{d}M$ (47a)
$\text{d}(M^{-1}M^{\prime})$ $=-\operatorname{Ad}_{M^{\prime-1}M}\text{d}M+\text{d}M^{\prime},$ (47b)
with $\operatorname{ad}$ the small adjoint on the Lie algebra. By the chain rule we obtain
$\text{d}(\operatorname{Ad}_{M^{-1}M^{\prime}}x)$ $=\operatorname{ad}_{\operatorname{Ad}_{M^{-1}M^{\prime}}x}\text{d}M-\operatorname{Ad}_{M^{-1}M^{\prime}}\operatorname{ad}_{x}\text{d}M^{\prime}.$ (48)
Applied to $M={}^{0}M_{c}(\bm{q})$ and $M^{\prime}={}^{0}M_{1}(\bm{q})$ we obtain
$\text{d}({}^{c}X_{1}(\bm{q})J_{1}v)$ $=\operatorname{ad}_{{}^{c}X_{1}J_{1}\bm{v}}\text{d}{}^{0}M_{c}-{}^{c}X_{1}\operatorname{ad}_{J_{1}\bm{v}}\text{d}{}^{0}M_{1}$ (49)
$=({}^{c}X_{1}J_{1}\bm{v})\times\text{d}{}^{0}M_{c}-{}^{c}X_{1}(J_{1}\bm{v}\times\text{d}{}^{0}M_{1}).$ (50)
Where in the second form we use the generalized cross product from the notation of Featherstone [16]. The calculus is similar for $J_{2}$. Now differentiating the collision detection, we have
$\frac{\text{d}{}^{0}M_{c}}{\text{d}\theta}$ $=\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{1}}J_{1}\frac{\text{d}\bm{q}}{\text{d}\theta}+\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{2}}J_{2}\frac{\text{d}\bm{q}}{\text{d}\theta},$ (51)
which can be computed using the randomized smoothed derivatives presented in Differentiable collision detection: a randomized smoothing approach [16] in the main paper. For self explanation of the paper, we provide the final formula: {mdframed}
$\frac{\partial J_{c}\bm{v}^{+}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ $=E\left[\left(({J_{c}\bm{v}^{+}})_{\times}\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{1}}-({}^{c}X_{1}{J_{1}\bm{v}^{+}})_{\times}\right)J_{1}+\left(({J_{c}\bm{v}^{+}})_{\times}\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{2}}+({}^{c}X_{2}{J_{2}\bm{v}^{+}})_{\times}\right)J_{2}\right]\frac{\text{d}\bm{q}}{\text{d}\theta}.$
To compute the $\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ we use the previous term and the duality stating that for any $\bm{\lambda}$ and $\bm{v}$ we have $\langle J_{c}^{\top}\bm{\lambda},\bm{v}\rangle=\langle\bm{\lambda},J_{c}\bm{v}\rangle$. Taking derivatives we have
$\langle\partial(J_{c}^{\top}\bm{\lambda})\text{d}\bm{q},\bm{v}\rangle$ $=\langle\lambda,\partial(J_{c}\bm{v})\text{d}\bm{q}\rangle$
$=\langle\bm{\lambda},L\bm{v}\rangle$
$=\langle L^{\top}\bm{\lambda},\bm{v}\rangle,$ (52)
and because it is true for any $\bm{v}$ we have $\partial(J_{c}^{\top}\bm{\lambda})\text{d}\bm{q}=L^{\top}\bm{\lambda}$. We calculate $L$ using the anti-commutativity of $\operatorname{ad}$:
$\partial(J_{c}\bm{v})\text{d}\bm{q}$ $=E\left[\left(\operatorname{ad}_{J_{c}\bm{v}}\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{1}}-{}^{c}X_{1}\operatorname{ad}_{J_{1}\bm{v}}\right)J_{1}+\left(\operatorname{ad}_{J_{c}\bm{v}}\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{2}}+{}^{c}X_{2}\operatorname{ad}_{J_{2}\bm{v}}\right)J_{2}\right]\text{d}\bm{q}$
$L\bm{v}$ $=-E\left[\left(\operatorname{ad}_{\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{1}}J_{1}\text{d}\bm{q}}J_{c}-{}^{c}X_{1}\operatorname{ad}_{J_{1}\text{d}\bm{q}}J_{1}\right)+\left(\operatorname{ad}_{\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{2}}\text{d}\bm{q}}J_{c}+{}^{c}X_{2}\operatorname{ad}_{J_{2}\text{d}\bm{q}}J_{2}\right)\right]\bm{v}.$ (53)
Taking the transpose and introducing the operator $P$ is the variable commutation of $\operatorname{ad}^{\top}$. Precisely for all $x$ in the Lie algebra and $y$ in the dual Lie algebra: $\operatorname{ad}^{\top}_{x}y=\operatorname{P}_{y}x$ we obtain:
$L^{\top}\bm{\lambda}$ $=-\left[\left(J_{c}^{\top}\operatorname{ad}_{\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{1}}J_{1}\text{d}\bm{q}}^{\top}-J_{1}^{\top}\operatorname{ad}^{\top}_{J_{1}\text{d}\bm{q}}{}^{c}X_{1}^{\top}\right)+\left(J_{c}^{\top}\operatorname{ad}_{\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{2}}J_{2}\text{d}\bm{q}}^{\top}+J_{2}^{\top}\operatorname{ad}^{\top}_{J_{2}\text{d}\bm{q}}{}^{c}X_{2}^{\top}\right)\right]E\bm{\lambda}$
$\partial(J_{c}^{\top}\bm{\lambda})\text{d}\bm{q}$ $=-\left[\left(J_{c}^{\top}\operatorname{P}_{E\bm{\lambda}}\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{1}}J_{1}-J_{1}^{\top}\operatorname{P}_{{}^{c}X_{1}^{\top}E\bm{\lambda}}J_{1}\right)+\left(J_{c}^{\top}\operatorname{P}_{E\bm{\lambda}}\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{2}}J_{2}+J_{2}^{\top}\operatorname{P}_{{}^{c}X_{2}^{\top}E\bm{\lambda}}J_{2}\right)\right]\text{d}\bm{q}.$ (54)
And finally, we obtain the second term {mdframed}
$\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ $=\left[\left(J_{1}^{\top}\operatorname{P}_{{}^{c}X_{1}^{\top}E\bm{\lambda}^{*}}-J_{c}^{\top}\operatorname{P}_{E\bm{\lambda}^{*}}\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{1}}\right)J_{1}-\left(J_{2}^{\top}\operatorname{P}_{{}^{c}X_{2}^{\top}E\bm{\lambda}^{*}}+J_{c}^{\top}\operatorname{P}_{E\bm{\lambda}^{*}}\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{2}}\right)J_{2}\right]\frac{\text{d}\bm{q}}{\text{d}\theta}.$
Details on implementation. Here, the terms are calculated for one contact. For multiple contacts, $\frac{\partial J_{c}\bm{v}^{+}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ is the concatenation of the terms for individual contacts and $\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ is the sum of the terms from each contacts. Note that for elements of dual spatial algebra $y=[f,m]$ we have the closed form $\operatorname{P}_{y}=\begin{pmatrix}0&f_{\times}\\ f_{\times}&m_{\times}\end{pmatrix}$. Note also that similarly to the kinematic Jacobians, the two terms can be computed efficiently by exploiting the sparsity induced by the kinematic structure.
Appendix D Complete simulation gradients expression
Combining the terms from the collision detection, the term from the velocity forward kinematic, and the main calculation presented in Section III, the complete expression of the simulation step derivative is {mdframed}
$\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}$ $=\left.\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}\right|_{\lambda=\lambda^{*}}-\Delta tM^{-1}J_{c}^{\top}CA^{-1}B\left(J_{c}\left.\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}\right|_{\lambda=\lambda^{*}}+\left.\frac{\text{d}J_{c}\bm{v}^{+}}{\text{d}\theta}\right|_{v=v^{+}}+\frac{\partial J_{c}\bm{v}^{+}}{\partial c}\frac{\text{d}c}{\text{d}\theta}\right),$
with
$\left.\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}\right|_{\lambda=\lambda^{*}}$ $=\frac{\text{d}\bm{v}}{\text{d}\theta}+\Delta t\left(\frac{\partial\text{UFD}}{\partial\bm{q}}\frac{\text{d}\bm{q}}{\text{d}\theta}+\frac{\partial\text{UFD}}{\partial\bm{v}}\frac{\text{d}\bm{v}}{\text{d}\theta}+\frac{\partial\text{UFD}}{\partial\bm{\tau}}\frac{\text{d}\bm{\tau}}{\text{d}\theta}+M^{-1}\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}\frac{\text{d}c}{\text{d}\theta}\right),$ (55)
$\left.\frac{\text{d}J_{c}\bm{v}^{+}}{\text{d}\theta}\right|_{v=v^{+}}$ $=\frac{\partial\text{FKV}(\bm{q},\bm{v}^{+},c)}{\partial\bm{q}}\frac{\text{d}\bm{q}}{\text{d}\theta},$ (56)
where $\text{FKV}(\bm{q},\bm{v}^{+},c)$ is the forward kinematic velocity that gives the velocity of the origin of the frame $c$ when the system is in configuration $\bm{q}$ with joint velocity $\bm{v}^{+}$. $A$, $B$, $C$ are as presented in Appendix B and $\frac{\partial J_{c}\bm{v}^{+}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$, $\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ are as presented in Appendix C.
Details on implementation. The partial derivatives $\frac{\partial\text{UFD}}{\partial\bm{q},\bm{v},\bm{\tau}}$, $\frac{\partial\text{FKV}}{\partial\bm{q}}$ and the term $M^{-1}(\bm{q})J_{c}^{\top}(\bm{q})$ can be efficiently computed via rigid-body algorithm as implemented in Pinocchio. The terms $\frac{\partial J_{c}\bm{v}^{+}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$, $\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$, $A$, $B$ and $C$ can be computed efficiently jointly with the ABA derivatives during forward and backward search of the kinematic tree to exploit its sparsity.
Baumgarte stabilization is often used in practice to prevent penetration errors from growing. The correction is integrated by adding terms to the expression of $g$
$g=J_{c}v_{f}+\frac{\Phi(q)}{\Delta t}-K_{p}\left[\frac{\Phi(q)}{\Delta t}\right]_{-}-K_{d}J_{c}v$ (57)
where $K_{p}$ and $K_{d}$ are the gains of the corrector. In the case of sticking or sliding contacts we have $(G\bm{\lambda}+g)_{N}=0$ and expanding the expression of $g$ yields
$(J_{c}v^{+})_{N}=K_{d}(J_{c}v)_{N}-(1-K_{p})\frac{\Phi(q)}{\Delta t}.$ (58)
Therefore, using a Baumgarte correction affects the derivative of the simulation. In particular, the derivatives of the proportional term involve $\Phi$ and thus should be handled when computing the derivatives of the collision detection. Differentiating the derivative term $K_{d}J_{c}v$ is done similarly to the $J_{c}v^{+}$ term i.e. via the Forward Kinematics derivatives. In more details, the term in parentheses of (D) becomes
$\frac{\text{d}G}{\text{d}\theta}+\frac{\text{d}g}{\text{d}\theta}=J_{c}\left.\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}\right|_{\lambda=\lambda^{*}}+\left.\frac{\text{d}J_{c}\bm{v}^{+}}{\text{d}\theta}\right|_{v=v^{+}}+\frac{\partial J_{c}\bm{v}^{+}}{\partial c}\frac{\text{d}c}{\text{d}\theta}+\frac{(1-K_{p})}{\Delta t}\frac{\text{d}\Phi(q)}{\text{d}\theta}-K_{d}J_{c}\frac{\text{d}\bm{v}}{\text{d}\theta}-K_{d}\left.\frac{\text{d}J_{c}\bm{v}}{\text{d}\theta}\right|_{v=v}$ (59)
Appendix E Additional experimental support
Figure 6: The differentiable simulator is used to find a control torque stabilizing various robotics systems: a Unitree Go1 in a standing position with a 10kg mass on its back (Left) or in a ”hand-stand” pose (Center) and a humanoid Unitree H1 in a ”push-up” pose (Right).
This section provides several additional experiments using our differentiable simulator to solve contact inverse dynamics problems on underactuated robotics systems. The considered problems, depicted in Fig. 6, are the following:
  • A Unitree Go1 is stabilized in a standing position with a 10kg mass put on its back;
  • A Unitree Go1 is stabilized in a ”hand-stand” pose;
  • A Unitree H1 humanoid is stabilized in a ”push-up” pose.
In every case, the robots are stabilized by optimizing the torque on the actuators. We refer to the video attached to this paper for more visualization of the experiments.
Appendix F Policy training details
CartPole Swing Up task involves stabilizing an underactuated pendulum in an upright position starting from the pendulum hanging downwards. The system comprises a 5-dimensional observation space (cart position $x$ nad velocity $\dot{x}$, pole angle $[\sin(\theta),\cos(\theta)]$ and angular velocity $\dot{\theta}$) and a 1-dimensional action space controlling the cart’s prismatic joint torque. The cart joint is constrained to $[-2\text{m},2\text{m}]$. and the reward function is defined, similar to [48], as:
$R=-\theta^{2}-0.1\dot{\theta}^{2}-0.05x^{2}-0.01\dot{x}^{2}$ (60)
Episodes run for 240 time steps without early termination, with randomly sampled initial states.
Hopper environment evaluates locomotion control of a single-legged robot evolving in a plane. The state space consists of 11 dimensions: base height, rotation, linear velocity (2D) angular velocity; joint angles (3D) and velocities (3D). The action space is 3-dimensional, controlling joint torques for the thigh, leg and foot. The smooth reward function proposed by [48] combines multiple objectives:
$R=R_{\text{velocity}}+R_{\text{height}}+R_{\text{posture}}-0.1\|a\|^{2}$ (61)
where $R_{\text{velocity}}=v_{x}$, and:
$R_{\text{height}}$ =$=\begin{cases}-200\Delta_{h}^{2},&\text{if }\Delta_{h}<0\\ \Delta_{h},&\text{if }\Delta_{h}\geq 0\end{cases},$ (62)
$\Delta_{h}$ $=\text{clip}(h+0.3,-1,0.3)$
$R_{\text{posture}}=1-\left(\frac{\theta}{30^{\circ}}\right)^{2}$ (63)
Episodes terminate after 1000 time steps or if the robot height falls below $-0.45\text{m}$.
PPO hyperparameters
TABLE IV: PPO Hyperparameters
Environment Horizon Length Parallel Envs. Minibatch Size
[…]
Hopper 32 256 1024
common to both environments include: $\gamma=0.99$ and $\lambda=0.95$ for Generalized Advantage Estimation (GAE) calculation, a learning rate of $43e-$ 4 for both actor and critic using $10$ mini batch epochs. Environment-specific settings are provided in Table IV.
SHAC hyperparameters are consistent across both environments: $16$ environments in parallel with a short horizon of $20$, $\gamma=0.99$ and $\lambda=0.95$ for GAE calculation, policy and critic learning rates $43e-$ 4, $16$ training iterations for critic with $4$ minibatches and target value network $\alpha=99$.